Multiply
(x/3 + 3/4) (3x/4 - 3/5)

Answer by stanbon(75887)

Multiply
(x/3 + 3/4) (3x/4 - 3/5)
---------------------
Use FOIL:
F: (x/3)(3x/4) = 3x^2/12 = x^2/4
O: (x/3)(-3/5) = -1/5x
I: (3/4)(3x/4) = 9x/16
L: (3/4)(-3/5) = -9/20
==========================
Combine to get:
x^2/4 + (9/16-1/5)x -9/20
x^2/4 + (29/80)x - 9/20
=======================
Cheers,
Stan H.